Evaluate the integral "\oint" ez/z2+1 dz where c is the circle |z|=3.


1
Expert's answer
2021-09-06T14:56:13-0400

"\int\limits_{|z| = 3} {\frac{{{e^z}}}{{{z^2} + 1}}dz = } \int\limits_{|z| = 3} {\frac{{{e^z}}}{{(z + i)(z - i)}}dz}"

Singular points of the integrand:

"{z_1} = i,\,\,{z_2} = - i" - simple poles.

Both points fall inside the circle "|z | = 3" :





Let's find the residues at these points:

"\mathop {res}\limits_{z = i} f(z) = \mathop {\lim }\limits_{z \to i} \frac{{{e^z}}}{{(z + i)(z - i)}}(z - i) = \mathop {\lim }\limits_{z \to i} \frac{{{e^z}}}{{z + i}} = \frac{{{e^i}}}{{2i}}"

"\mathop {res}\limits_{z = - i} f(z) = \mathop {\lim }\limits_{z \to i} \frac{{{e^z}}}{{(z + i)(z - i)}}(z + i) = \mathop {\lim }\limits_{z \to i} \frac{{{e^z}}}{{z - i}} = \frac{{{e^{ - i}}}}{{ - 2i}}"

Then

"\int\limits_{|z| = 3} {\frac{{{e^z}}}{{{z^2} + 1}}dz = } = 2\pi i\sum {\mathop {res}\limits_{z = {z_k}} } f(z) = 2\pi i\left( {\frac{{{e^i}}}{{2i}} - \frac{{{e^{ - i}}}}{{2i}}} \right) = \pi \left( {{e^i} - {e^{ - i}}} \right)"

Answer:"\pi \left( {{e^i} - {e^{ - i}}} \right)"